### v1

# AeroGPT: Leveraging Large-Scale Audio Model for Aero-Engine Bearing Fault Diagnosis

Abstract
Aerospace engines, as critical components in aviation and aerospace industries, require continuous and accurate fault diagnosis to ensure operational safety and prevent catastrophic failures. While deep learning techniques have been extensively studied in this context, they output logits or confidence scores, necessitating post-processing to derive actionable insights. Furthermore, the potential of large-scale audio models in this domain remains largely untapped. To address these limitations, this paper proposes AeroGPT, a novel framework that transfers knowledge from general audio domain to aero-engine bearing fault diagnosis. AeroGPT is a framework based on large-scale audio model that incorporates Vibration Signal Alignment (VSA) to adapt general audio knowledge to domain-specific vibration patterns, and combines Generative Fault Classification (GFC) to directly output interpretable fault labels. This approach eliminates the need for post-processing of fault labels, supports interactive, interpretable, and actionable fault diagnosis, thereby greatly enhancing industrial applicability. Through comprehensive experimental validation on two aero-engine bearing datasets, AeroGPT achieved exceptional performance with 98.94% accuracy on the DIRG dataset and perfect 100% classification on the HIT bearing dataset, surpassing traditional deep learning approaches. Additional Qualitative analysis validates the effectiveness of our approach and highlights the potential of large-scale models to revolutionize fault diagnosis.
Index Terms:
Aerospace Engine, Deep Learning, Large Language Model, Fault Diagnosis
I Introduction
Aerospace engines serve as the cornerstone of modern aviation and aerospace industries, powering everything from commercial aircraft to space exploration vehicles [1, 2]. Their reliable operation is paramount to aviation safety, mission success, and the prevention of catastrophic failures that could result in significant human and economic losses. Within these sophisticated propulsion systems, aero-engine bearings are particularly critical components, operating under extreme conditions of high speeds, temperature variations, and significant mechanical stress [3, 4, 5]. The failure of these bearings can lead to severe engine malfunctions, making them a primary focus for condition monitoring and maintenance protocols [6, 2]. Fault diagnosis of aero-engine bearings has consequently emerged as an essential discipline, offering methodologies to detect, identify, and predict potential failures before they manifest as catastrophic events. Early and accurate bearing fault detection can significantly extend engine lifespan, optimize maintenance schedules, and ensure operational safety [7]. Traditionally, bearing fault diagnosis approaches rely on a two-step process of feature extraction followed by classification [8]. For instance, Medina et al. [9] proposed an approach using Poincare plots for feature extraction, with subsequent classification performed by a multi-class Support Vector Machine (SVM). However, such approaches are limited by their dependence on domain-specific feature engineering and conventional machine learning classifiers, which may fail to fully capture the complex dynamics of aero-engine bearing systems.
In recent years, data-driven methods, particularly deep learning approaches, have revolutionized fault diagnosis by enabling end-to-end learning directly from raw vibration signals [10, 11, 12]. These methods automatically extract hierarchical features, eliminating the need for manual feature engineering that traditional approaches require. Convolutional Neural Networks (CNNs) have been widely adopted for this purpose, as they excel at capturing local patterns and spatial hierarchies in data. Wang et al. [13] proposed a one-dimensional CNN with attention mechanism tailored for bearing fault enhancement and classification, establishing a new state-of-the-art performance on the wheelset bearing dataset. Apart from CNNs, Recurrent Neural Networks, especially Long Short-Term Memory (LSTM) networks, have also gained traction in vibration signal analysis. LSTMs are adept at capturing temporal dependencies in sequential data, making them suitable for modeling the time-varying nature of vibration signals. Song et al. [14] developed an CNN-BiLSTM network that combines CNN’s spatial feature extraction capabilities with BiLSTM’s temporal modeling, demonstrating robust performance even with limited samples. Transformer-based architectures [15], leveraging self-attention mechanisms to capture long-range dependencies and contextual information, have also emerged as powerful tools for vibration signal analysis. For instance, Li et al. [16] proposed Dconformer, a novel CNN-Transformer network that combines joint-learning denoising with a multi-branch cross-cascaded architecture to extract both local and global features from noisy vibration signals. Xiang et al. [17] introduced a frequency channel-attention based Vision Transformer method, which integrates frequency domain characteristics with self-attention mechanisms to enhance model interpretability while maintaining computational efficiency through a shift linear layer design. Alongside architectural innovations, significant research has focused on learning strategies to enhance model adaptability and reduce data dependency. Prominent examples include transfer learning, which adapts models pre-trained on large source domains [18, 19]; meta-learning [20] and the closely related field of few-shot learning [21, 22], which aim to train models capable of recognizing new fault types from very limited data; and unsupervised learning [23], which leverages large amounts of unlabeled data.
However, the data-driven methods mentioned above typically output logits or confidence scores, necessitating post-processing to derive actionable insights, whereas industrial applications require efficient and interpretable solutions. Recently, the emrgency and development of large language models (LLMs) and large multimodal models (LMMs) offer transformative advantages in this regard. These models have demonstrated remarkable capabilities in generating human-like text, understanding complex queries, and providing contextually relevant responses. With interactivity, they can incorporate contextual inputs or user queries to provide tailored, situation-specific diagnoses. Furthermore, they can be designed to directly output decisions or actionable recommendations, eliminating the need for complex post-processing. More critically, LLMs shift the paradigm from low-information-density fault categories to high-information-density analysis, delivering detailed insights more than just defect categories, thereby enhancing both interpretability and practical utility in industrial applications.
Despite these advances, the application of large-scale models to fault diagnosis remains in its early stages, with several noteworthy contributions emerging recently. BearingFM [24] established a framework for training foundation models in this domain, utilizing domain knowledge-based data augmentation and contrastive learning to extract features from unlabeled vibration signals, achieving high accuracy with minimal labeled data. Tao et al. [25] leveraged the text modality by quantitatively selecting features of vibration signals to textualize the time series data and utilized Low-Rank Adaptation (LoRA) for fine-tuning LLMs, demonstrating how language models can interpret bearing conditions. Taking a different approach, FaultGPT [26] exploited the vision modality by extracting features from vibration time-frequency images, which were then paired with textual instructions to generate detailed diagnostic reports.
Building upon these observations, we identify a critical oversight in current approaches: whether utilizing the text modality of LLMs or extending to the visual modality of LMMs, these methods fail to recognize a fundamental characteristic of bearing vibration signals, which is their intrinsic similarity to acoustic phenomena. Bearing vibration signals, as mechanical oscillations propagating through a medium, manifest as time-varying waveforms with frequency, amplitude, and temporal patterns that closely resemble those of sound waves. This acoustic likeness is evident in their shared spectral properties, such as harmonic peaks and resonance frequencies, which are routinely analyzed in audio processing but underutilized in fault diagnosis. Furthermore, despite the critical importance of aerospace propulsion systems and the severe consequences of their failure, large-scale models specifically designed for aero-engine bearing fault diagnosis have yet to be developed. The absence of such specialized models represents a significant gap in the current research landscape and limits the effectiveness of predictive maintenance in aerospace applications.
Motivated by the above challenges, this paper aims to align general domain audio knowledge with aero-engine bearing vibration patterns, and for aerospace engine bearing fault diagnosis, proposes a novel framework, AeroGPT. As a large-scale audio model-based approach, AeroGPT addresses the limitations of current methods through two key innovations: Vibration Signal Alignment (VSA), which adapts general audio knowledge to the specific characteristics of aero-engine bearing vibrations, and Generative Fault Classification (GFC), which retains the generative capabilities of the underlying LLM to directly output interpretable fault labels, eliminating the need for post-processing. By leveraging the inherent generative capabilities of large-scale audio models, AeroGPT enables interactive, interpretable, and actionable fault diagnosis, greatly enhancing its practical utility in aerospace applications. Our experimental validation on an aerospace bearing dataset demonstrates that AeroGPT achieves an impressive 98.94% accuracy on the DIRG dataset and 100% classification accuracy on the HIT bearing dataset, outperforming traditional deep learning approaches. Additional qualitative analyses further validate the effectiveness of our approach to support interactive fault diagnosis and follow-up analysis, showcasing the interpretability and actionability of the generated outputs. These results underscore the potential of large-scale audio models to revolutionize fault diagnosis in critical aerospace systems.
The main contributions of this paper can be summarized as follows:
  1. 1.
    To the best of our knowledge, this paper is the first to adapt large-scale audio model for industrial fault diagnosis, proposing a new paradigm that differs from the deep learning techniques previously used in this field.
  2. 2.
    This paper proposes Vibration Signal Alignment (VSA), which bridges the gap between general audio knowledge and domain-specific vibration patterns, enabling the model to adapt its acoustic representations to the specific characteristics of bearing vibration signal.
  3. 3.
    This paper proposes Generative Fault Classification (GFC), leveraging the inherent generative capabilities of LLMs to directly output interpretable, actionable insights rather than traditional logits that require post-processing.
  4. 4.
    This paper proposes AeroGPT, a novel framework based on large-scale audio model that transfers the knowledge in general audio domain to aero-engine bearing fault diagnosis while retaining the generative capabilities of the underlying LLM. Experimental results demonstrate that AeroGPT achieves state-of-the-art performance and highlight the potential of large-scale models to revolutionize fault diagnosis.
The remainder of this paper is structured as follows: Section II introduces the methodology of the proposed AeroGPT framework, including the overview of the system architecture, audio knowledge acquisition process, LoRA-based domain knowledge adaptation techniques, Vibration Signal Alignment (VSA) mechanism, and Generative Fault Classification (GFC) approach; Section III presents the experimental validation, including comprehensive evaluations on the DIRG and HIT bearing datasets; Section IV concludes the paper by summarizing the contributions and findings.
II Methodology
II-A Overview of AeroGPT
Figure 1: Demonstration of the overall framework of AeroGPT, along with its practical application in the aero-engine bearing fault diagnosis process.
Figure 2: Technical components of the AeroGPT methodology. The framework is initialized with a foundation model, followed by Vibration Signal Alignment (VSA) to adapt general audio knowledge to domain-specific vibration patterns and Vibration Signal Alignment (GFC) to output interpretable fault labels.
For aero-engine bearing fault diagnosis, this paper proposes AeroGPT, a novel framework leveraging large-scale audio models, and the overall pipeline is shown in Fig. 1. This framework explores for the first time the application of large audio models to aerospace bearing fault diagnosis, with the aim of generating interpretable and actionable diagnostic outputs directly from vibration signals. AeroGPT transforms the conventional classification-based fault diagnosis approach into an interactive, generative process capable of providing detailed insights beyond simple fault categorization.
As shown in Fig. 1, the vibration signals of the aero-engine bearings are collected from the bearing test rig and then input into an audio encoder equipped with a LoRA layer, which effectively converts the vibration patterns into audio tokens for LLM understanding. The input representation process can be formalized as:
$\mathcal{I}_{\text{AeroGPT}}=\{\mathcal{E}_{\text{audio}}(\mathbf{v}_{\text{proc}}),\mathcal{E}_{\text{text}}(\mathbf{p}),\mathcal{E}_{\text{text}}(\mathbf{c})\}$ (1)
where $\mathcal{E}_{\text{audio}}$ and $\mathcal{E}_{\text{text}}$ are the audio and text encoders respectively, $\mathbf{v}_{\text{proc}}$ is the processed vibration signal, $\mathbf{p}$ is the prompt template, and $\mathbf{c}$ represents optional context information. The audio tokens and text tokens are then concatenated together as the input for the LLM, which then processes these tokens and generates a fault label as the output. Leveraging the generative capabilities of the LLM, analysis and follow-up questions can also be conducted interactively, allowing for a more comprehensive understanding of the fault condition.
AeroGPT incorporates two key mechanisms that address fundamental challenges in utilizing large-scale audio models for fault diagnosis, as illustrated in Fig. 2. The first component, Vibration Signal Alignment (VSA), bridges the gap between general audio knowledge and the specific characteristics of aero-engine bearing vibrations. VSA processes paired inputs of vibration signals and corresponding textual descriptions, enabling the model to establish meaningful connections between acoustic patterns and their diagnostic interpretations. The second one, Generative Fault Classification (GFC), leverages the inherent capabilities of LLMs to directly generate diagnostic outputs in natural language rather than numerical logits that require post-processing. This approach provides actionable insights for maintenance personnel and supports interactive diagnostic sessions where follow-up questions can be addressed, demonstrating great practical utility of the system in aerospace applications. Detailed descriptions of the components are provided in the following sections.
II-B Audio Knowledge Acquisition
The acoustic understanding capabilities in this framework are developed through a comprehensive multi-stage process, as illustrated in the leftmost part of Fig. 2. This process systematically builds robust audio pattern recognition abilities that can be transferred to bearing vibration signal analysis.
The initial stage of knowledge acquisition involves exposure to diverse audio-related tasks. As shown in the diagram, this includes transcription tasks where audio inputs (represented by the microphone icon) are converted into textual outputs (e.g., “[en]Hello there!”). This process enables the model to develop fundamental representations of acoustic patterns and their corresponding semantic interpretations. By processing varied audio inputs through structured transcription tasks, the model learns to extract meaningful features from complex acoustic signals.
Building upon basic acoustic understanding, the second stage employs supervised fine-tuning (SFT) to enhance the model’s analytical capabilities. This stage focuses on higher-level audio understanding tasks, in which the model learns to address complex queries about audio content, such as “What does the speaker feel?” and “What is the sound?”, providing responses like “Very happy.”, “Sound of rain.” This stage cultivates the model’s ability to extract contextual information and identify subtle acoustic features. The model also learns to understand human intent through instruction pairs.
The final refinement stage implements a Direct Preference Optimization (DPO) approach. This technique uses a comparative scoring mechanism where alternative responses to the same audio input are evaluated, with scores indicating their relative quality (e.g., Answer 1 with score 8.5 vs. Answer 2 with score 2.0). The optimization objective is formulated as:
$\mathcal{L}_{DPO}=-\mathbb{E}\Biggl{[}\log\sigma\Biggl{(}\beta\log\frac{\mathcal{P}_{\theta}(y_{w}|x)}{\mathcal{P}_{ref}(y_{w}|x)}-\beta\log\frac{\mathcal{P}_{\theta}(y_{l}|x)}{\mathcal{P}_{ref}(y_{l}|x)}\Biggr{)}\Biggr{]},$ (2)
where $\mathcal{P}_{\theta}$ represents the system being optimized, $\mathcal{P}_{ref}$ is a reference system, $y_{w}$ and $y_{l}$ are the preferred and less preferred responses respectively, $\sigma$ is the sigmoid function, and $\beta$ is a hyperparameter controlling optimization strength. This comparative approach ensures the system’s acoustic interpretations align with expert assessments, particularly for challenging or ambiguous acoustic patterns.
The culmination of these three stages produces a large-scale audio foundation model with sophisticated acoustic pattern recognition capabilities. This audio knowledge provides the foundation for AeroGPT’s ability to interpret vibration signals from aero-engine bearings as acoustic phenomena, enabling more accurate and nuanced fault diagnosis than conventional approaches.
II-C LoRA-based Domain Knowledge Adaptation
To effectively transfer general audio knowledge to aero-engine bearing fault diagnosis, AeroGPT employs a domain adaptation approach based on Low-Rank Adaptation (LoRA) techniques. This adaptation enables efficient knowledge transfer while maintaining computational feasibility even with limited resources.
The adaptation process operates on the principle that domain-specific knowledge can be effectively integrated through low-rank decomposition of weight updates. Rather than modifying the entire weight matrices of the audio processing system, which would require enormous computational resources, AeroGPT freezes the pre-trained weight matrices and introduces trainable low-rank adaptations. For a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ in the model, the adaptation is accomplished through:
$\mathbf{W}=\mathbf{W}_{0}+\Delta\mathbf{W}=\mathbf{W}_{0}+\mathbf{B}\mathbf{A}$ (3)
where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$ are low-rank matrices with rank $r\ll\min(d,k)$. This factorization significantly reduces the number of trainable parameters from $d\times k$ to $r\times(d+k)$.
During the adaptation process, the original pre-trained weight matrix $\mathbf{W}_{0}$ remains frozen, while only the low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ are updated. The forward pass computation is modified as:
$\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\Delta\mathbf{W}\mathbf{x}=\mathbf{W}_{0}\mathbf{x}+\mathbf{B}\mathbf{A}\mathbf{x}$ (4)
where $\mathbf{x}$ represents the input to the layer and $\mathbf{h}$ is the output. To facilitate stable training, matrix $\mathbf{A}$ is typically initialized with random Gaussian values, while matrix $\mathbf{B}$ is initialized with zeros, ensuring that at the beginning of training, $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}=0$.
In practice, a scaling factor $\alpha$ is also introduced to control the magnitude of the adaptation:
$\mathbf{h}=\mathbf{W}_{0}\mathbf{x}+\frac{\alpha}{r}\mathbf{B}\mathbf{A}\mathbf{x}$ (5)
Here, $\alpha$ is a hyperparameter that can be set equal to the rank $r$ or tuned separately. This scaling factor helps stabilize training, especially when experimenting with different rank values.
The LoRA adaptation technique employed in AeroGPT offers unparalleled parameter efficiency with over 99% parameter reduction compared to full fine-tuning approaches. This efficiency translates to significant reductions in VRAM usage, from over 100GB for full fine-tuning to approximately 17GB with LoRA, making the fine-tuning process feasible on widely available hardware. This is particularly advantageous in aerospace applications, where computational resources may be limited.
In our framework, LoRA-based adaption is strategically applied at multiple points. Within the audio encoder, LoRA layers modify the processing of vibration signals, enabling the system to recognize the unique acoustic signatures of various bearing faults. For the language model component, adaptation points are established at the self-attention mechanisms (specifically in query, key, and value projections) and feed-forward networks. Through this knowledge adaptation process, AeroGPT is able to bridge the gap between general audio understanding and the specific characteristics of aero-engine bearing vibrations, establishing a robust foundation for effective fault diagnosis.
II-D Vibration Signal Alignment (VSA)
The Vibration Signal Alignment (VSA) mechanism is designed to bridge the fundamental gap between general audio knowledge and the domain-specific characteristics of aero-engine bearing vibrations. This is achieved through a alignment process that preserves the rich hierarchical representations inherent in large-scale audio models while adapting them to the specialized patterns of mechanical vibrations.
The proper alignment requires establishing connections between vibration signals and their corresponding semantic interpretations. To facilitate this, we construct a dataset of vibration signals paired with detailed textual descriptions that articulate their diagnostic significance. Each description systematically captures critical fault-related information, including vibration source, noise characteristics, fault type, severity, location, and relevant operating conditions. This pairing strategy enables the model to develop a nuanced understanding of the relationship between acoustic features and their diagnostic implications. To enhance generalizability and prevent overfitting, the vibration data contains diverse bearing sources that differ from those used in our final aerospace validation tests.
The alignment process begins with precise signal preparation to ensure optimal encoding. Raw vibration signals $\mathbf{v}_{\text{raw}}\in\mathbb{R}^{T}$ undergo amplitude normalization through a statistical transformation:
$\mathbf{v}_{\text{proc}}=\mathcal{F}_{\text{norm}}(\mathbf{v}_{\text{raw}};\alpha,\beta)=\beta\cdot\frac{\mathbf{v}_{\text{raw}}-\mu_{\mathbf{v}}}{\sigma_{\mathbf{v}}}+\alpha$ (6)
where $\mu_{\mathbf{v}}$ and $\sigma_{\mathbf{v}}$ are the mean and standard deviation of the signal, and $\alpha,\beta$ are calibration hyperparameters optimized for the audio encoder’s dynamic range. This normalization ensures consistent amplitude profiles across diverse operational conditions and sensor configurations, significantly enhancing the model’s robustness to variations in signal acquisition parameters.
The normalized vibration signals are then transformed into audio embeddings through our domain-adapted encoder:
$\mathbf{z}_{\text{audio}}=\mathcal{E}_{\text{audio}}(\mathbf{v}_{\text{proc}};\theta_{\text{base}}+\Delta\theta_{\text{LoRA}})$ (7)
where $\mathcal{E}_{\text{audio}}$ represents the audio encoder, $\theta_{\text{base}}$ denotes the frozen parameters of the pre-trained model, and $\Delta\theta_{\text{LoRA}}$ encompasses the trainable LoRA parameters. This formulation enables selective adaptation of the encoder’s parameters while preserving its foundational acoustic understanding. The encoder projects temporal vibration patterns into a high-dimensional embedding space $\mathbf{z}_{\text{audio}}\in\mathbb{R}^{L_{a}\times d}$, where $L_{a}$ represents the sequence length and $d$ is the embedding dimension.
In parallel, textual prompts and context information are processed through a text encoder to obtain corresponding embeddings:
$\mathbf{z}_{\text{text}}=\mathcal{E}_{\text{text}}(\mathbf{p},\mathbf{c};\theta_{\text{text}})\in\mathbb{R}^{L_{t}\times d}$ (8)
where $\mathbf{p}$ represents the instruction prompt, $\mathbf{c}$ denotes optional context, $\theta_{\text{text}}$ are the parameters of the text encoder, $L_{t}$ is the text sequence length, and $d$ is the embedding dimension shared with the audio encoder to facilitate cross-modal interactions.
These audio embeddings are subsequently tokenized to create a discrete representation compatible with the language model’s input format:
$\mathbf{T}_{\text{audio}}=\text{Tokenize}(\mathbf{z}_{\text{audio}})=\{\mathbf{t}^{(a)}_{1},\mathbf{t}^{(a)}_{2},...,\mathbf{t}^{(a)}_{L_{a}}\}\in\mathbb{R}^{L_{a}\times d}$ (9)
where $\mathbf{T}_{\text{audio}}$ represents the sequence of audio tokens of length $L_{a}$, each with dimension $d$, derived from the encoded vibration signal.
The cross-modal attention mechanism then establishes connections between the audio and text modalities, computing attention weights $\mathbf{A}\in\mathbb{R}^{L_{t}\times L_{a}}$ as:
$\mathbf{A}=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)=\text{softmax}\left(\frac{\mathbf{z}_{\text{text}}\mathbf{W}_{Q}(\mathbf{z}_{\text{audio}}\mathbf{W}_{K})^{\top}}{\sqrt{d_{k}}}\right)$ (10)
where $\mathbf{W}_{Q}$ and $\mathbf{W}_{K}$ are learnable projection matrices, and $d_{k}$ is the dimensionality of the key vectors. This mechanism allows the model to selectively attend to relevant regions of the vibration signal when interpreting or generating diagnostic text, effectively establishing a vibration-to-language mapping that captures subtle fault-indicative patterns.
The training objective is to predict the next token in an autoregressive manner, with the main goal being to optimize the model so that it can predict subsequent tokens in the text description based on vibration signals and previous text. During training process, the cross-entropy loss is minimized:
$\mathcal{L}_{\text{CE}}=-\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{v},\mathbf{y})\in\mathcal{D}}\sum_{t=1}^{|\mathbf{y}|}\log p_{\theta}(y_{t}|\mathbf{z}_{\text{audio}},y_{<t})$ (11)
where $\mathcal{D}$ represents the training dataset of vibration-text pairs, $\mathbf{y}$ denotes the target textual description, $y_{t}$ is the $t$-th token in the description, $y_{<t}$ represents all preceding tokens, and $p_{\theta}$ is the probability distribution over the vocabulary predicted by the model with parameters $\theta$.
This training paradigm enables the model to learn contextual relationships between vibration patterns and diagnostic interpretations in an end-to-end manner while preserving the linguistic capabilities of the underlying large language model, allowing for the generation interpretable diagnostic outputs rather than mere classification logits. Our experimental results demonstrate that this alignment strategy is able to transfer the general audio knowledge to the specific domain of aero-engine bearing vibrations, enhancing the model’s ability to recognize and interpret complex fault patterns.
II-E Generative Fault Classification (GFC)
The Generative Fault Classification (GFC) represents a paradigm shift in fault diagnosis by leveraging the inherent generative capabilities of large language models to directly produce interpretable diagnostic outputs. Unlike conventional deep learning approaches that output numerical logits requiring post-processing, GFC enables AeroGPT to generate human-readable fault labels and explanations in natural language, significantly enhancing both interpretability and actionability for maintenance personnel.
Traditional fault classification methodologies typically employ multi-class classifiers that output probability distributions across predefined fault categories, necessitating argmax operations and label mapping to determine the final classification. In contrast, GFC capitalizes on the autoregressive nature of LLMs to generate textual fault labels directly. This approach eliminates intermediate post-processing steps, streamlining the diagnostic pipeline while providing outputs in a format that is immediately comprehensible to maintenance technicians. The generation process follows:
$\hat{y}_{t}=\underset{y\in\mathcal{V}}{\arg\max}\>p_{\theta}(y|\mathbf{z}_{\text{audio}},\hat{y}_{<t})$ (12)
where $\hat{y}_{t}$ represents the predicted token at position $t$, $\mathcal{V}$ denotes the vocabulary space, $\mathbf{z}_{\text{audio}}$ is the encoded representation of the vibration signal, and $\hat{y}_{<t}$ comprises all previously generated tokens. This autoregressive generation continues until a complete fault diagnosis is produced.
For model optimization, the training objective is formulated as a cross-entropy loss function over the sequence of tokens constituting the fault label:
$\mathcal{L}_{\text{GFC}}=-\frac{1}{|\mathcal{D}_{\text{fault}}|}\sum_{(\mathbf{v},\mathbf{y})\in\mathcal{D}_{\text{fault}}}\sum_{t=1}^{|\mathbf{y}|}\log p_{\theta}(y_{t}|\mathbf{z}_{\text{audio}},y_{<t})$ (13)
where $\mathcal{D}_{\text{fault}}$ represents the fault diagnosis dataset comprising vibration signal-label pairs $(\mathbf{v},\mathbf{y})$, and $p_{\theta}(y_{t}|\mathbf{z}_{\text{audio}},y_{<t})$ denotes the probability of generating the correct token $y_{t}$ given the audio embedding and preceding tokens. During parameter optimization, we utilize LoRA to selectively adapt the model while preserving its foundational generative capabilities, which can be represented by:
$\theta_{\text{LoRA}}^{*}=\underset{\theta_{\text{LoRA}}}{\arg\min}\>\mathbb{E}_{(\mathbf{v},\mathbf{y})\sim\mathcal{D}_{\text{fault}}}\left[-\log p_{\theta_{\text{base}}\oplus\theta_{\text{LoRA}}}(\mathbf{y}|\mathbf{v})\right]$ (14)
where $\theta_{\text{base}}\oplus\theta_{\text{LoRA}}$ denotes the composition of frozen base parameters with trainable LoRA parameters, enabling efficient adaptation while mitigating catastrophic forgetting of the model’s generative capabilities.
A distinctive advantage of GFC is its seamless integration with follow-up analysis capabilities. By preserving the generative nature of the underlying LLM, AeroGPT can not only classify faults but also provide interpretable labels and respond to queries about fault characteristics. This is formalized as a conditional generation task:
$p_{\theta}(\mathbf{r}|\mathbf{v},\mathbf{y},\mathbf{q})=\prod_{t=1}^{|\mathbf{r}|}p_{\theta}(r_{t}|\mathbf{z}_{\text{audio}},\mathbf{y},\mathbf{q},r_{<t})$ (15)
where $\mathbf{r}$ represents the generated response, $\mathbf{q}$ denotes a follow-up query, and $\mathbf{y}$ is the initially generated fault label. This process enables contextual analysis that considers both the original vibration signal and the diagnostic history, facilitating deeper exploration of fault characteristics and implications.
GFC offers several substantial advantages over traditional classification approaches. It eliminates post-processing steps such as argmax operations and label mapping, which streamlines the diagnostic pipeline. The natural language outputs enhance interpretability by being immediately comprehensible to maintenance personnel without requiring specialized knowledge of model architectures or output interpretations. Additionally, the generative nature of GFC enables straightforward extensibility to new fault types without requiring architectural modifications, as the system can adapt to evolving aerospace industry simply by including examples of new faults in the training data. Experimental results demonstrate that GFC not only achieves superior classification accuracy compared to traditional approaches but also provides richer diagnostic information with practical utility for maintenance operations in aerospace applications.
III Experimental Validation
III-A Experimental Setup
To systematically evaluate the performance of AeroGPT, a series of experiments were conducted using two widely-recognized bearing datasets: the aerospace bearing dataset from DIRG and the aero-engine bearing dataset from HIT. These datasets represent diverse operational conditions and fault characteristics relevant to aerospace applications, providing a robust foundation for validating our approach.
The experimental validation encompasses three primary parts: First, an ablation study to assess the individual contributions of each component within the AeroGPT framework, specifically examining the impact of the Foundation Model (FM), Vibration Signal Alignment (VSA), and Generative Fault Classification (GFC). Second, comparative evaluations against state-of-the-art deep learning models to benchmark AeroGPT’s performance in terms of accuracy, precision, and F1-score. Thrid, qualitative analyses to demonstrate AeroGPT’s capability to generate interpretable, text-based diagnostic outputs that provide insights beyond mere fault classifications.
III-A1 Evaluation Metrics
To quantitatively assess AeroGPT’s diagnostic performance, we employed standard classification metrics including accuracy, precision, and F1-score. These metrics provide complementary perspectives on the model’s effectiveness in correctly identifying bearing faults across diverse operational conditions:
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (16)
$\text{Precision}=\frac{TP}{TP+FP}$ (17)
$\text{Recall}=\frac{TP}{TP+FN}$ (18)
[…]
Here, TP (True Positive) represents correctly identified faulty conditions, TN (True Negative) denotes correctly identified healthy conditions, FP (False Positive) indicates healthy conditions incorrectly classified as faulty, and FN (False Negative) represents faulty conditions incorrectly classified as healthy. Additionally, we evaluated the model’s performance separately for non-defective and defective categories to assess its balanced capability across different fault conditions.
Beyond these quantitative metrics, the qualitative aspects of AeroGPT’s outputs were also evaluated, particularly focusing on their interactivity, interpretability, and actionability. This dual evaluation approach provides a comprehensive assessment of both classification accuracy and practical utility in aerospace maintenance contexts.
III-A2 Implementation Details
The implementation of AeroGPT was developed using the PyTorch framework, with the HuggingFace Transformers library utilized for model architecture and training utilities. All experiments were conducted on a single NVIDIA A10 GPU with 24GB of VRAM.
The foundation model used in the experiments was initialized on the weights of Qwen2-Audio [27], which comprises a pre-trained audio encoder with 124 million parameters and a large language model with 7 billion parameters. To fit the input of audio encoder, the vibration signals were sampled at 16kHz and segmented into wav files. For the Vibration Signal Alignment (VSA) stage, a dataset of 36,053 vibration-text paired samples was utilized to bridge the domain gap between general audio knowledge and bearing-specific vibration patterns. The datasets used for classification were split into training and test sets, with a split ratio of 8:2.
During fine-tuning process, the LoRA rank was set to 16, with a scaling factor of 32. The learning rate was initialized at $1\times 10^{-5}$, and a linear warm-up schedule was employed for the first 5% of total training steps to ensure stability. The batch size was set to 32, with gradient accumulation over 16 steps to maximize GPU utilization while maintaining memory efficiency.
III-B Case 1: DIRG Bearing Dataset
III-B1 Dataset Description
The primary dataset for experimental validation was obtained from the Dynamic and Identification Research Group (DIRG) at Politecnico di Torino [28]. This aerospace-focused dataset features high-speed aeronautical bearings operating at speeds up to 35,000 rpm, accurately representing aerospace application conditions. Their bearing test rig comprises a shaft supported by three roller bearings, with controlled fault conditions introduced to one of them. Vibration data was collected using triaxial IEPE accelerometers mounted at two strategic locations on the bearing supports. The dataset encompasses seven distinct bearing conditions: one healthy state (0A) and six damage states (1A-6A) with precisely controlled conical indentations (150, 250, and 450 $\mu$m) on either the inner ring or a single roller. Data was recorded across multiple operational scenarios with shaft speeds ranging from 6,000 to 30,000 rpm and radial loads varying from 0 to 1800 N, at a sampling frequency of 51,200 Hz. To simulate the harsh operational environment, 4 dB of Gaussian white noise was added to the original signals. A total number of 22,134 training samples and 7,259 test samples were utilized.
TABLE I: Ablation study results demonstrating the contributions of each component in AeroGPT.
FM VSA GFC Accuracy Precision F1-score
$\checkmark$ 14.87% 6.31% 4.20%
$\checkmark$ $\checkmark$ 20.65% 12.47% 10.83%
$\checkmark$ $\checkmark$ 97.21% 97.84% 97.52%
$\checkmark$ $\checkmark$ $\checkmark$ 98.94% 99.16% 99.02%
III-B2 Ablation Study
To comprehensively evaluate the contributions of each component in the proposed AeroGPT framework, we conducted an ablation study. TABLE I presents the performance metrics when including different combinations of the three primary components: Foundation Model (FM), which represents the model which has gone through Audio Knowledge Acquisition process, with general audio understanding capabilities; Vibration Signal Alignment (VSA), which bridges the domain gap between general audio and aero-engine bearing signal domain; and Generative Fault Classification (GFC), which employs fine-tuning techniques for direct fault label generation. All experiments maintained identical testing conditions with the same dataset split and evaluation metrics to ensure fair comparisons.
The results reveal that utilizing only the Foundation Model (FM) without any adaptation yields poor performance, with merely 14.87% accuracy, 6.31% precision, and 4.20% F1-score, which is approximately equivalent to random guessing. This underscores the significant domain gap between general audio understanding and aerospace bearing vibration analysis. When incorporating the VSA component while omitting GFC, a modest improvement can be observed in performance—accuracy increases by 5.78% to 20.65%, precision nearly doubles to 12.47%, and F1-score improves by 6.63% to 10.83%. This enhancement, while limited, indicates that the VSA stage establishes preliminary connections between general audio knowledge and vibration-specific patterns, though insufficient for accurate fault classification in isolation.
Notably, integrating GFC directly with the FM without the VSA intermediary yields a dramatic performance leap. The FM+GFC configuration achieves 97.21% accuracy, 97.84% precision and 97.52% F1-score, representing absolute improvements of 82.34%, 91.53%, and 93.32% in the respective metrics compared to using FM alone. This substantial enhancement demonstrates that PEFT enables effective adaptation of the foundation model to the target task, even without explicit domain alignment. The generative approach to fault classification proves highly effective in capturing the discriminative patterns in aerospace bearing vibration signals.
The complete AeroGPT framework, incorporating all three components (FM+VSA+GFC), achieves the highest performance across all metrics: 98.94% accuracy, 99.16% precision, and 99.02% F1-score. Compared to the FM+GFC configuration, the full framework yields further absolute improvements of 1.73%, 1.32%, and 1.50% in accuracy, precision, and F1-score, respectively. This corresponds to relative error reductions of 62.1% in misclassification rate, 61.1% in precision errors, and 60.5% in F1-score deficiency. These results confirm the synergistic effect of the VSA stage, which, while providing minimal benefits in isolation, significantly enhances the effectiveness of the subsequent GFC stage by aligning the foundation model’s representations with the aerospace bearing vibration signal domain.
The significant gap between configurations with and without GFC highlights the critical role of task-specific fine-tuning in adapting general-purpose models to specialized domains like aerospace bearing fault diagnosis. However, the performance improvement from adding VSA to GFC demonstrates that bridging the domain gap through intermediate pre-training remains valuable, especially for highly specialized technical applications such as aerospace propulsion systems monitoring.
III-B3 Comparison with Existing Methods
TABLE II: Performance comparison of AeroGPT against traditional deep learning models on the DIRG dataset.
Method Non-Defective Defective Total
Accuracy Precision F1-score Accuracy Precision F1-score Accuracy Precision F1-score
AeroGPT 99.46% 97.08% 98.14% 98.42% 99.26% 99.07% 98.94% 99.16% 99.02%
MA1DCNN 99.15% 99.18% 96.96% 93.26% 95.73% 96.08% 96.21% 96.31% 96.22%
ResNet50 98.14% 93.42% 93.47% 87.35% 92.63% 92.63% 92.75% 92.97% 92.78%
ConvNeXt V2 95.55% 84.53% 84.40% 72.05% 83.68% 83.70% 83.80% 83.77% 83.76%
Conv-Transformer 96.11% 86.27% 86.40% 71.16% 83.19% 83.17% 83.64% 83.77% 83.32%
Masked AutoEncoder 98.29% 94.62% 93.98% 84.96% 91.13% 91.23% 91.62% 91.61% 91.61%
PoolFormer 94.79% 80.94% 82.02% 67.19% 81.00% 80.82% 80.99% 80.99% 80.97%
Swin-Transformer 95.91% 87.83% 85.26% 79.27% 87.55% 87.96% 87.59% 87.82% 87.65%
Swin-Transformer V2 97.60% 90.59% 91.71% 80.69% 88.90% 88.71% 89.14% 89.27% 89.19%
To comprehensively evaluate the efficacy of the proposed AeroGPT framework, we conducted comparative experiments against a diverse range of advanced deep learning architectures representing the current state-of-the-art in deep-learning-based classification tasks. The benchmark models span multiple architectural paradigms, including: MA1DCNN [13], a specialized multi-scale attention-enhanced convolutional architecture optimized for bearing vibration analysis; ResNet50 [29], a deep residual network that mitigates gradient vanishing through skip connections; ConvNeXt V2 [30], a modern convolutional design incorporating depthwise convolutions and layer normalization; Conv-Transformer, a hybrid architecture leveraging both local convolutional feature extraction and global attention mechanisms; Masked AutoEncoder [31], a self-supervised learning approach that reconstructs masked input segments; PoolFormer [32], which replaces attention mechanisms with simpler pooling operations while maintaining competitive performance; and the hierarchical Swin-Transformer architectures (original [33] and V2 [34]), which compute self-attention within shifted windows to achieve linear computational complexity with respect to image size. For vision-based architectures like ResNet and Swin-Transformer, vibration signals were converted to two-dimensional feature maps before processing.
As evidenced in TABLE II, AeroGPT achieves substantial performance advantages across nearly all evaluation metrics, with particularly pronounced improvements in defective state recognition. The overall accuracy of 98.94% represents a 2.73% improvement over the next best performer (MA1DCNN at 96.21%) and more than 8 percentage points above most transformer-based architectures. This performance gap underscores a fundamental limitation in conventional approaches: despite their sophisticated architectures, models designed for generic feature extraction struggle to capture the patterns of bearing faults without domain-specific adaptation.
A distinctive pattern emerges when examining performance across defective and non-defective categories. Traditional architectures exhibit a substantial performance asymmetry, with accuracy discrepancies exceeding 20 percentage points in several cases (e.g., ConvNeXt V2: 95.55% vs. 72.05%; Conv-Transformer: 96.11% vs. 71.16%; PoolFormer: 94.79% vs. 67.19%). This indicates a fundamental challenge in discriminating between different fault types compared to distinguishing between healthy and faulty states. In contrast, AeroGPT demonstrates remarkable category balance, with comparable performance across both non-defective (99.46%) and defective (98.42%) classifications. This balanced capability is particularly critical in aerospace applications, where missed defects can lead to catastrophic failures and false alarms result in unnecessary and costly maintenance interventions.
The performance advantage of AeroGPT stems from its architectural advantages. First, unlike conventional models that build representations from scratch, AeroGPT leverages and adapts the rich acoustic domain knowledge, which provides a sophisticated basis for interpreting mechanical vibration patterns that share fundamental properties with acoustic phenomena. Second, the VSA stage effectively bridges the domain gap between general audio and aerospace-specific vibration patterns, as confirmed by the ablation study results.
Notably, even the most advanced transformer-based architectures like Swin-Transformer V2, which achieve state-of-the-art performance in many computer vision tasks, fall significantly short in this domain-specific application. This emphasizes that architectural sophistication alone is insufficient for specialized engineering tasks—domain knowledge integration.
III-B4 Qualitative Analysis
Figure 3: Examples of AeroGPT’s generative fault diagnosis capability and answers to follow-up queries.
Beyond quantitative metrics, AeroGPT’s interactive diagnostic capabilities were evaluated to assess practical utility in aerospace maintenance scenarios. Fig. 3 demonstrates how AeroGPT leverages its generative foundation to transform fault diagnosis from isolated classification to interactive consultation, enabling multifaceted analysis of bearing defects.
Instead of producing a simple fault label, AeroGPT generates a detailed fault description when initially queried about defect presence. The framework’s output explicitly identifies the fault as “inner ring indentation” with a severity of “150 mm,” providing immediate context for maintenance personnel. When subsequently questioned about severity, the system demonstrates nuanced assessment capabilities, classifying the defect as “moderate” while noting it “should be monitored closely to prevent further deterioration.” This level of granular evaluation is not typically available in traditional classification outputs, which often lack interpretability and actionable insights.
Moreover, AeroGPT exhibits causal reasoning by establishing explicit connections between vibration patterns and their mechanical implications. When asked about performance impacts, it articulates how inner ring indentation directly causes “increased vibration” and reduced “efficiency and lifespan” of the bearing component. This mechanical insight contrasts sharply with traditional classifiers that provide no explanation of fault mechanisms or consequences, merely indicating fault presence and category.
The framework further demonstrates domain expertise through contextually appropriate technical recommendations. When queried about inspection methodologies, AeroGPT recommends specific diagnostic tools, “a borescope for visual inspection and an ultrasonic device to measure the depth of the indentation”, which are tailored to the identified fault type. Similarly, when asked about factors that might exacerbate the condition, the system identifies “high-speed rotation, heavy loads, or high temperatures” as operational parameters that could accelerate deterioration, providing valuable risk management insights for maintenance scheduling.
AeroGPT’s ability to generate contextually appropriate maintenance recommendations calibrated to aerospace standards represents perhaps its most significant operational advantage. The response prescribing a “maintenance check to inspect the bearing and potentially replace the inner ring” aligns with established protocols for aerospace bearing defects of this category and severity. Furthermore, AeroGPT can incorporate domain-specific contextual information from technical maintenance manuals as additional input, enabling fine-grained recommendation calibration that precisely adheres to manufacturer specifications or organizational maintenance procedures, further enhancing recommendation precision. This multifaceted capability effectively bridges the critical gap between fault detection and maintenance action that typically requires expert human interpretation of classification outputs.
These examples confirm that AeroGPT’s generative approach fundamentally transforms fault diagnosis methodology through a comprehensive set of capabilities spanning detection, assessment, causal analysis, technical guidance, risk evaluation, and intervention planning—all delivered through natural language interaction. The system’s ability to reason about mechanical implications and prescribe appropriate interventions represents a paradigm shift in the interaction between maintenance personnel and diagnostic systems in safety-critical aerospace applications. By providing this complete diagnostic workflow within a conversational interface, AeroGPT significantly reduces the expertise barrier for maintenance personnel, potentially accelerating response times and improving maintenance outcomes in aerospace operations.
III-C Case 2: HIT Bearing Dataset
III-C1 Dataset Descrption
The second dataset utilized was obtained by Hou et al. from the Harbin Institute of Technology (HIT) [35], featuring inter-shaft bearing data from an aero-engine system. Their experimental setup comprises a modified aero-engine with the critical dual-rotor structure preserved. The test rig follows a similar structure with that in the first case, with the low-pressure rotor connected to the high-pressure rotor through a shaft and the inter-shaft bearings subjecting to various fault conditions. The system is equipped with two independent motors for the high and low pressure rotors, allowing for precise control over operational parameters. All signals were sampled at 25kHz, with tests conducted across 28 distinct operating conditions varying both speed (ranging from 1000 r/min to 6000 r/min) and speed ratio between the rotors (1.2 to 1.8). The dataset encompasses 2412 sets of vibration data labeled into three categories (healthy, inner ring defect, and outer ring defect), segmented into 20480-point sequences. The obtained training set consists of 1929 samples and the test set contains 483 samples.
III-C2 Comparison with Existing Methods
TABLE III: Performance comparison of AeroGPT against traditional deep learning models on the HIT dataset.
Methods Accuracy Precision F1-score
AeroGPT 100.00% 100.00% 100.00%
MA1DCNN 99.59% 99.59% 99.59%
ResNet50 99.79% 99.79% 99.79%
ConvNeXt V2 94.41% 95.71% 95.94%
Conv-Transformer 99.17% 99.18% 99.17%
Masked AutoEncoder 98.34% 98.47% 98.34%
PoolFormer 99.59% 99.50% 99.75%
Swin-Transformer 99.80% 99.75% 99.88 %
Swin-Transformer V2 99.59% 99.50% 99.49%
Experimental results on the HIT inter-shaft bearing dataset demonstrate that AeroGPT achieves perfect classification performance with 100.00% accuracy, precision, and F1-score across both defective and non-defective categories, as shown in TABLE III. Comparative models also performed exceptionally well on this dataset. While the performance gap between AeroGPT and traditional approaches is narrower on this dataset compared to DIRG, AeroGPT maintains its advantage by eliminating all misclassifications. This consistent superiority across both aerospace bearing datasets confirms the effectiveness of transferring audio domain knowledge to aero-engine bearing fault diagnosis through the proposed methodologies.
III-C3 Qualitative Analysis
Figure 4: Comparison of AeroGPT’s generative fault diagnosis ability with general-purpose models and conventional approaches.
The comparative qualitative evaluation illustrated in Fig. 4 demonstrates AeroGPT’s distinctive advantages over both general-purpose large language models and traditional deep learning approaches when diagnosing aero-engine bearing faults.
When presented with identical vibration signals, AeroGPT produces a definitive diagnosis with high specificity, stating unequivocally “Yes, there is an indentation on the inner ring.” This precise identification contrasts sharply with the responses from other large language models, which exhibit significant diagnostic uncertainty despite similar architectural capabilities. The large foundation audio model, Qwen2-Audio-Instruct, produces the tentative assessment “Yes, the signal indicates that there may be a defect,” while the general-purpose model ChatGPT similarly suggests ‘Yes, the yellow part indicates that there may be a defect in the bearing.” The hedging language reflects these models’ lack of specialized knowledge alignment with bearing vibration patterns, resulting in diagnostically ambiguous conclusions that offer limited practical utility for maintenance personnel.
The diagnostic specificity gap becomes even more pronounced when users request detailed fault information. When asked to distinguish between inner and outer ring defects, AeroGPT maintains its diagnostic precision, confidently identifying the inner ring location. In contrast, both Qwen2-Audio and ChatGPT acknowledge a fundamental limitation with the response “There is no sufficient information to determine.” This inability to provide location-specific diagnosis represents a significant practical limitation, as repair protocols and maintenance prioritization depend heavily on defect location within the bearing assembly.
Traditional deep learning methods present an entirely different set of limitations. While potentially achieving comparable classification accuracy, these approaches output numerical probability distributions (e.g., [0.13, 0.82, 0.05]) that require additional post-processing and expert interpretation to translate into actionable diagnoses. The necessity of this translation layer introduces potential inconsistency and significantly limits practical utility in maintenance contexts. The workflow typically involves programming and training a model, obtaining a probability distribution across classes, and then performing post-processing to map the highest probability (0.82 in this case) to a predefined class label (Class 1). Even after this multi-step process, the system remains inherently non-interactive, with the notation “unable to interact with” highlighting a fundamental limitation—these systems cannot engage in follow-up diagnostic discussions or provide explanations for their classifications.
AeroGPT eliminates these limitations through its knowledge-aligned generative approach. By directly outputting precise fault characterizations in natural language and supporting location-specific diagnoses without additional post-processing, it bridges the interpretability gap that typically separates diagnostic systems from maintenance personnel. This capability transforms the diagnostic paradigm from a complex multi-stage analysis requiring specialized expertise to an intuitive conversational process accessible to broader maintenance staff. The directness and specificity of AeroGPT’s diagnostic outputs represent a significant advancement in practical utility for aerospace applications, where maintenance decisions must be made promptly and with high confidence to ensure operational safety.
IV Conclusion
This paper proposed AeroGPT, a novel framework based on large-scale audio model that transfers knowledge from general audio domains to aero-engine bearing fault diagnosis. By recognizing the intrinsic acoustic-like nature of bearing vibration signals, AeroGPT addresses fundamental limitations in current aero-engine fault diagnosis approaches. The framework’s two key innovations, Vibration Signal Alignment (VSA) and Generative Fault Classification (GFC), provide a systematic methodology for adapting general audio knowledge to domain-specific vibration patterns while enabling direct generation of interpretable diagnostic outputs. Through comprehensive experimental validation on two aerospace bearing datasets, AeroGPT achieved exceptional performance with 98.94% accuracy on the DIRG dataset and perfect 100% classification on the HIT bearing dataset, surpassing traditional deep learning approaches. The qualitative analysis further demonstrated AeroGPT’s unique capability to provide definitive, specific fault characterizations in natural language, contrasting sharply with the uncertainty exhibited by general-purpose language models and the post-processing requirements of conventional classification approaches. This transformation of fault diagnosis from complex analysis to intuitive conversation significantly enhances practical utility in aerospace maintenance contexts, where rapid, accurate interpretation is essential for preventing catastrophic failures. By eliminating the need for post-processing and providing interactive, interpretable, and actionable diagnostics, AeroGPT represents a significant advancement in aerospace reliability engineering, highlighting the substantial potential of large-scale audio models to revolutionize fault diagnosis in industrial settings.